OpenStudy (_fluffeh_):

WILL MEDAL What is the product in lowest terms? 5 / 6 times - 8 / 15 A. -8/9 B. -4/9 C. -1/3 D. -1/7

OpenStudy (_fluffeh_):

@undeadknight26

OpenStudy (anonymous):

B

OpenStudy (_fluffeh_):

CoC says you need to explain, please explain >_<

OpenStudy (anonymous):

5/6 × -8/15 = (5×-8) / (6×15) = -40/90 = -4/9

OpenStudy (anonymous):

THAT WHAT I GOT :)

OpenStudy (_fluffeh_):

Thank you! ^u^

OpenStudy (anonymous):

AND IF YOU can search up fraction calculator it helped me and it HAS THE EXPLNATION

undeadknight26 (undeadknight26):

Yup what they said